# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@ 2.3. Fourier–Mukai transform
Define $\tilde{X}$ to be the GIT quotient
$\tilde{X}=\tilde{V}\sslash_{\tilde{\theta}}\tilde{G}.$
The morphism $f_{-}:\tilde{X}\to X_{-}$ is induced by the morphism of stacks
$[\tilde{V}/\tilde{G}]\to[V/G]$
defined by the map
$\tilde{V}$ $\to V$
$(X,U,Y)$ $\mapsto(X,YU)$
@@
$\tilde{G}$ $\to G$
$(g_{1},g_{2})$ $\mapsto g_{1}.$
There is a similar description of the map $f_{+}:\tilde{X}\to X_{+}.$ The $\mathbb{T}$-action on $E$ and $F$ induces an action on $\tilde{X}$ for which the maps $f_{\pm}:\tilde{X}\to X_{\pm}$ are equivariant. We define
(2.3.1) $\mathbb{FM}:K^{0}_{\mathbb{T}}(X_{-})\to K^{0}_{\mathbb{T}}(X_{+})$
@@
In [BLVdB16], it is proven that the associated functor $\mathbb{FM}:D^{b}(X_{-})\to D^{b}(X_{+})$, between derived categories is an equivalence in the case that $B$ is a point. In [BCF+21] it is shown that this equivalence may also be described using windows. The arguments naturally extend to our setting. We will use this fact only to conclude that the map (2.3.1) is an isomorphism and preserves the pairing.
We compute $\mathbb{FM}$ using localization as in [Tho92]. The fixed loci of $\tilde{X}$ are indexed by a pair of size-$r$ subsets of $\{1,\ldots,n\}$, corresponding to choosing $r\times r$ minors of both $\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ and $\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$, i.e. they are indexed by elements of $D_{-}\times D_{+}$.
Let $R_{1},R_{2}$ be the representations of $\tilde{G}$ obtained by composing the projection $\pi_{1},\pi_{2}:\tilde{G}\to G$ to $G$ with the right representation $R$ of $G$. Again we abuse notation and also use $R_{1/2}$ to denote the induced vector bundles on $\tilde{X}$. With these conventions, we observe that $f_{-/+}^{*}(R)=R_{1/2}$.
@@
For $(\delta^{-},\delta^{+})\in D_{-}\times D_{+}$, let $B_{(\delta^{-},\delta^{+})}$ denote the associated fixed locus, isomorphic to $B$. Then $f_{+}$ (resp. $f_{-}$) maps $B_{(\delta^{-},\delta^{+})}$ isomorphically onto $B_{\delta^{+}}$ (resp. $B_{\delta^{-}}$). Furthermore, $(R_{1})|_{B_{(\delta^{-},\delta^{+})}}\cong\oplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}$ and $(R_{2})|_{B_{(\delta^{-},\delta^{+})}}\cong\oplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}$.
@@
$e_{\delta^{-}}=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\mathcal{H}om(M_{j_{-}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})^{\vee}$
(2.3.2) $\cong$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}$
in $K^{0}_{\mathbb{T}}(X_{-})$, where for $A\in K^{0}_{\mathbb{T}}(X_{-})$, we denote $\wedge^{\bullet}A:=\oplus_{i=0}^{\operatorname{rk}(A)}(-1)^{i}\wedge^{i}A$.
@@
The $\mathbb{T}$-localized equivariant $K$-theory of $X_{-}$ is spanned by the classes $\{\pi_{-}^{*}(A)\otimes e_{\delta^{-}}|A\in K^{0}(B),\delta^{-}\in D_{-}\}$.
@@
By Lemma 2.1, one checks that the class $\pi_{-}^{*}(A)\otimes e_{\delta^{-}}$ restricts to zero on $B_{\delta_{0}^{-}}$ for $\delta_{0}^{-}\neq\delta^{-}$. Furthermore, it restricts to the invertible class $\wedge^{\bullet}N_{B_{\delta^{-}}|\operatorname{Gr}(r,F^{\vee})}^{\vee}$ on $B_{\delta^{-}}$. The result then follows from the localization theorem in $K$-theory. ∎
@@
(2.3.3) $\mathbb{FM}(e_{\delta^{-}})=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}.$
@@
Note that although the right hand side of the equation above is the same formal expression as the right hand side of (2.3.2), the right hand side of (2.3.3) is to be interpreted as a class in $K^{0}_{\mathbb{T}}(X_{+})$.
@@
Note first that, in $K^{0}_{\mathbb{T}}(\tilde{X})$, the class
$\prod_{i_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{i_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}\cdot\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}$
restricts to $\wedge^{\bullet}N_{B_{\delta^{-},\delta^{+}}|\tilde{X}}^{\vee}$ on $B_{(\delta^{-},\delta^{+})}$, and restricts to zero on $B_{(\delta^{-}_{1},\delta^{+}_{1})}$ for $(\delta^{-}_{1},\delta^{+}_{1})\neq(\delta^{-},\delta^{+})$.
It follows that
$f_{-}^{*}(e_{\delta^{-}})=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}$
$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}}{\prod_{i_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{i_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}\cdot\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}}\right)$
$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{1}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}}\right)$
$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{2}^{\vee})^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}(M_{j}^{\vee}\otimes R_{2}^{\vee})^{\vee}}\right)$
where the second equality is by the localization theorem applied to $K^{0}_{\mathbb{T}}(\tilde{X})$, the fourth is Lemma 2.4. Pushing forward via $f_{+}$, we obtain
$(f_{+})_{*}f_{-}^{*}(e_{\delta^{-}})=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{+}})_{*}i_{\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R)^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}(M_{j}^{\vee}\otimes R^{\vee})^{\vee}}\right)$
$=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}$
where the last equality is the localization theorem applied to $K^{0}_{\mathbb{T}}(X_{+})$.
@@
(2.3.4) $C_{\delta^{-},\delta^{+}}=$ $\prod_{i=1}^{r}e^{(n-r)(x_{\delta^{+}_{i}}-z_{\delta^{-}_{i}})/2}\prod_{j_{-}\notin\delta^{-}}\frac{\sin((x_{\delta^{+}_{i}}-z_{j_{-}})/2\mathbbm{i})}{\sin((z_{\delta^{-}_{i}}-z_{j_{-}})/2\mathbbm{i})},$
where $\mathbbm{i}=\sqrt{-1}$.
@@
Define $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to be the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{\delta^{-}})\in H^{*}_{\mathbb{T}}(B_{\delta^{-}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to
(2.3.5) $\sum_{\delta^{+}\in D_{+}}C_{\delta^{-},\delta^{+}}\frac{\phi_{\delta^{+},\delta^{-}}(\alpha)}{e_{\mathbb{T}}(N_{\delta^{+}})}$
where $\phi_{\delta^{+},\delta^{-}}:H^{*}_{\mathbb{T}}(B_{\delta^{-}})\to H^{*}_{\mathbb{T}}(B_{\delta^{+}})$ is the canonical isomorphism induced by the projections $\pi_{\delta^{+}}:B_{\delta^{+}}\to B$ and $\pi_{\delta^{-}}:B_{\delta^{-}}\to B$.
@@
We will prove $\operatorname{ch}^{\mathbb{T}}(\mathbb{FM}(e_{\delta^{-}}))=\mathbb{U}_{H}(\operatorname{ch}^{\mathbb{T}}(e_{\delta^{-}}))$. The general statement then follows by Corollary 2.5 and the projection formula. Observe that
$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{-}}^{*}e_{\delta^{-}})=$ $\operatorname{ch}^{\mathbb{T}}\left(i_{\delta^{-}}^{*}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}\right)$
$=$ $\operatorname{ch}^{\mathbb{T}}\left(\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes M_{\delta^{-}_{i}})^{\vee}\right)$
$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\left(1-e^{z_{\delta^{-}_{i}}-z_{j_{-}}}\right)$
@@
$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{+}}^{*}\mathbb{FM}(e_{\delta^{-}}))=$ $\operatorname{ch}^{\mathbb{T}}i_{\delta^{+}}^{*}\left(\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}\right)$
$=$ $\operatorname{ch}^{\mathbb{T}}\left(\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes L_{\delta^{+}_{i}})^{\vee}\right)$
$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\left(1-e^{x_{\delta^{+}_{i}}-z_{j_{-}}}\right)$
@@
$\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{\left(1-e^{x_{\delta^{+}_{i}}-z_{j_{-}}}\right)}{\left(1-e^{z_{\delta^{-}_{i}}-z_{j_{-}}}\right)}$
$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{e^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\left(e^{-(x_{\delta^{+}_{i}}-z_{j_{-}})/2}-e^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\right)}{e^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\left(e^{-(z_{\delta^{-}_{i}}-z_{j_{-}})/2}-e^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\right)}$
$=$ $C_{\delta^{-},\delta^{+}}.$
By the localization theorem in cohomology (see [AB84]), it follows that the diagram commutes. ∎
@@
${\pi_{-}}_{*}\left(\alpha\cup\operatorname{Td}_{X_{-}/B}\right)={\pi_{+}}_{*}\left(\mathbb{U}_{H}(\alpha)\cup\operatorname{Td}_{X_{+}/B}\right).$

### v2

@@ 2.3 Fourier–Mukai transform
Define $\tilde{X}$ to be the GIT quotient, $\tilde{X}=\tilde{V}\sslash_{\tilde{\theta}}\tilde{G}$. The morphism $f_{-}\colon\tilde{X}\to X_{-}$ is induced by the morphism of stacks $\bigl{[}\tilde{V}/\tilde{G}\bigr{]}\to[V/G]$ defined by the map
$\tilde{V}\to V,\qquad\bigl{(}\tilde{M},\tilde{D},\tilde{N}\bigr{)}\mapsto\bigl{(}\tilde{M},\tilde{N}\tilde{D}\bigr{)}$
@@
$\tilde{G}\to G,\qquad(g_{1},g_{2})\mapsto g_{1}.$
There is a similar description of the map $f_{+}\colon\tilde{X}\to X_{+}.$ The $\mathbb{T}$-action on $E$ and $F$ induces an action on $\tilde{X}$ for which the maps $f_{\pm}\colon\tilde{X}\to X_{\pm}$ are equivariant. We define
$\mathbb{FM}\colon\ K^{0}_{\mathbb{T}}(X_{-})\to K^{0}_{\mathbb{T}}(X_{+})$ (2.3)
@@
Lemma 2.4.
Let $H=\operatorname{tot}\bigl{(}F^{\vee}\otimes E\bigr{)}=\operatorname{tot}(\mathcal{H}{\rm om}(F,E))$. The space $\tilde{X}$ is equal to the fiber product $X_{+}\times_{H}X_{-}.$
Proof.
It suffices to show that $\tilde{X}$ satisfies the universal property of fiber products. Suppose that $q\colon S\to B$ is a $B$-scheme. A morphism $S\to X_{+}$ over $B$ corresponds to a rank $r$ vector bundle $U_{+}\to S$ together with linear maps $M_{+}\colon F\to U_{+}$ and $N_{+}\colon U_{+}\to E$ such that $N_{+}$ is full rank. A morphism $S\to X_{-}$ over $B$ corresponds to a rank $r$ vector bundle $U_{-}\to S$ together with linear maps $M_{-}\colon F\to U_{-}$, $N_{-}\colon U_{-}\to E$ such that $M_{-}$ is full rank.
On the other hand, a morphism $S\to\tilde{X}$ corresponds to a pair of rank $r$ vector bundles $U_{+/-}\to S$ (obtained via the induced map $S\to B{\rm GL}_{r}\times B{\rm GL}_{r}$) together with linear maps $\tilde{M}\colon F\to U_{-}$, $\tilde{D}\colon U_{-}\to U_{+}$, and $\tilde{N}\colon U_{+}\to E,$ such that the first and last are both full rank.
Suppose we have morphisms $S\to X_{+/-}$ which agree when composed with the natural maps $X_{+/-}\to H$. This yields two triples $(U_{+},M_{+},N_{+})$ and $(U_{-},M_{-},N_{-})$ such that $N_{+}\circ M_{+}=N_{-}\circ M_{-}$. We define a morphism $S\to\tilde{X}$ as follows. Note first that since $N_{+}$ is full rank, the condition $N_{+}\circ M_{+}=N_{-}\circ M_{-}$ implies that $N_{-}$ factors through a map $\tilde{D}\colon U_{-}\to U_{+}$. Then define $\tilde{M}=M_{-}$ and $\tilde{N}=N_{+}.$ The triple $\tilde{M}$, $\tilde{D}$, $\tilde{N}$ defines a map $S\to\tilde{X}$ as in the previous paragraph.
It is easy to check that the compositions $S\to\tilde{X}\to X_{+/-}$ agree with the original morphisms $S\to X_{+/-}$ from the previous paragraph. Thus $\tilde{X}$ satisfies the universal property as desired. ∎
In [9, Theorem D], it is proven that the associated functor $\mathbb{FM}\colon D^{b}(X_{-})\to D^{b}\bigl{(}X_{+}\bigr{)}$, between derived categories is an equivalence in the case that $B$ is a point. [^2] In [9], the Fourier–Mukai kernel used is $\mathcal{O}_{X_{+}\times_{H}X_{-}}$. This agrees with our kernel $\mathcal{O}_{\tilde{X}}$ by Lemma 2.4. In [3], it is shown that this equivalence may also be described using windows. The arguments naturally extend to our setting. We will use this fact only to conclude that the map (2.3) is an isomorphism and preserves the pairing.
We compute $\mathbb{FM}$ using localization as in [32]. The fixed loci of $\tilde{X}$ are indexed by a pair of size-$r$ subsets of $\{1,\dots,n\}$, corresponding to choosing $r\times r$ minors of both $\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ and $\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$, i.e., they are indexed by elements of $D_{-}\times D_{+}$.
Let $R_{1}$, $R_{2}$ be the representations of $\tilde{G}$ obtained by composing the projection $\pi_{1},\pi_{2}\colon\tilde{G}\to G$ to $G$ with the right representation $R$ of $G$. Again, we abuse notation and also use $R_{1/2}$ to denote the induced vector bundles on $\tilde{X}$. With these conventions, we observe that $f_{-/+}^{*}(R)=R_{1/2}$.
@@
For $(\delta^{-},\delta^{+})\in D_{-}\times D_{+}$, let $B_{(\delta^{-},\delta^{+})}$ denote the associated fixed locus, isomorphic to $B$. Then $f_{+}$ ((((resp. $f_{-})$ maps $B_{(\delta^{-},\delta^{+})}$ isomorphically onto $B_{\delta^{+}}$ ((((resp. $B_{\delta^{-}})$. Furthermore, $(R_{1})|_{B_{(\delta^{-},\delta^{+})}}\cong\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}$ and $(R_{2})|_{B_{(\delta^{-},\delta^{+})}}\cong\bigoplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}$.
@@
$e_{\delta^{-}}=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\mathcal{H}{\rm om}(M_{j_{-}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})^{\vee}\cong\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}$ (2.4)
in $K^{0}_{\mathbb{T}}(X_{-})$, where for $A\in K^{0}_{\mathbb{T}}(X_{-})$, we denote $\wedge^{\bullet}A:=\bigoplus_{i=0}^{\operatorname{rk}(A)}(-1)^{i}\wedge^{i}A$.
@@
The $\mathbb{T}$-localized equivariant $K$-theory of $X_{-}$ is spanned by the classes $\bigl{\{}\pi_{-}^{*}(A)\otimes e_{\delta^{-}}|A\in K^{0}(B),\delta^{-}\in D_{-}\bigr{\}}$.
@@
By Lemma 2.1, one checks that the class $\pi_{-}^{*}(A)\otimes e_{\delta^{-}}$ restricts to zero on $B_{\delta_{0}^{-}}$ for $\delta_{0}^{-}\neq\delta^{-}$. Furthermore, generalizing the argument of [12, Section 5.1] to the relative setting, we see that the relative tangent bundle of $\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}$ is obtained from the following relative Euler sequence for $\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}$:
$0\rightarrow S_{-}\otimes S_{-}^{\vee}\rightarrow S_{-}^{\vee}\otimes F^{\vee}\rightarrow T\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}/B\rightarrow 0$ (2.5)
Therefore, $\pi_{-}^{*}(A)\otimes e_{\delta^{-}}$ restricts to $A\otimes\wedge^{\bullet}N_{B_{\delta^{-}}|\operatorname{Gr}(r,F^{\vee})}^{\vee}$ where $\wedge^{\bullet}N_{B_{\delta^{-}}|\operatorname{Gr}(r,F^{\vee})}^{\vee}$ is an invertible class on $B_{\delta^{-}}$. The result then follows from the localization theorem in $K$-theory as in [30, Section 2.3]. ∎
@@
$\mathbb{FM}(e_{\delta^{-}})=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}.$ (2.6)
@@
Note that although the right-hand side of the equation above is the same formal expression as the right-hand side of (2.4), the right-hand side of (2.6) is to be interpreted as a class in $K^{0}_{\mathbb{T}}\bigl{(}X_{+}\bigr{)}$.
@@
Note first that, in $K^{0}_{\mathbb{T}}\bigl{(}\tilde{X}\bigr{)}$, the class
$\prod_{i_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{i_{-}}^{\vee}\otimes R_{1}^{\vee}\bigr{)}^{\vee}\cdot\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}\bigl{(}R_{2}^{\vee}\otimes R_{1}\bigr{)}^{\vee}$
restricts to $\wedge^{\bullet}N_{B_{\delta^{-},\delta^{+}}|\tilde{X}}^{\vee}$ on $B_{(\delta^{-},\delta^{+})}$, and restricts to zero on $B_{(\delta^{-}_{1},\delta^{+}_{1})}$ for $\bigl{(}\delta^{-}_{1},\delta^{+}_{1}\bigr{)}\neq(\delta^{-},\delta^{+})$. It follows that
$f_{-}^{*}(e_{\delta^{-}})=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R_{1}^{\vee}\bigr{)}^{\vee}$
${}=\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}$
[…]
${}=\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\Biggl{(}\frac{1}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}\bigl{(}R_{2}^{\vee}\otimes R_{1}\bigr{)}^{\vee}}\Biggr{)}$
${}=\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\Biggl{(}\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R_{2}^{\vee}\bigr{)}^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}\bigl{(}M_{j}^{\vee}\otimes R_{2}^{\vee}\bigr{)}^{\vee}}\Biggr{)},$
where the second equality is by the localization theorem applied to $K^{0}_{\mathbb{T}}\bigl{(}\tilde{X}\bigr{)}$, the fourth is Lemma 2.5. Pushing forward via $f_{+}$, we obtain
$(f_{+})_{*}f_{-}^{*}(e_{\delta^{-}})={}$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{+}})_{*}i_{\delta^{+}}^{*}\Biggl{(}\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R)^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}\bigl{(}M_{j}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}}\Biggr{)}$
[…]
where the last equality is the localization theorem applied to $K^{0}_{\mathbb{T}}(X_{+})$. ∎
@@
$C_{\delta^{-},\delta^{+}}={}$ $\prod_{i=1}^{r}{\rm e}^{(n-r)(x_{\delta^{+}_{i}}-z_{\delta^{-}_{i}})/2}\prod_{j_{-}\notin\delta^{-}}\frac{\sin\bigl{(}\bigl{(}x_{\delta^{+}_{i}}-z_{j_{-}}\bigr{)}/2{\rm i}\bigr{)}}{\sin\bigl{(}\bigl{(}z_{\delta^{-}_{i}}-z_{j_{-}}\bigr{)}/2{\rm i}\bigr{)}},$ (2.7)
where ${\rm i}=\sqrt{-1}$.
@@
Define $\mathbb{U}_{H}\colon H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to be the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{\delta^{-}})\in H^{*}_{\mathbb{T}}(B_{\delta^{-}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to
$\sum_{\delta^{+}\in D_{+}}C_{\delta^{-},\delta^{+}}\frac{\phi_{\delta^{+},\delta^{-}}(\alpha)}{e_{\mathbb{T}}(N_{\delta^{+}})},$ (2.8)
where $\phi_{\delta^{+},\delta^{-}}\colon H^{*}_{\mathbb{T}}(B_{\delta^{-}})\to H^{*}_{\mathbb{T}}(B_{\delta^{+}})$ is the canonical isomorphism induced by the projections $\pi_{\delta^{+}}\colon B_{\delta^{+}}\to B$ and $\pi_{\delta^{-}}\colon B_{\delta^{-}}\to B$.
@@
We will prove $\operatorname{ch}^{\mathbb{T}}(\mathbb{FM}(e_{\delta^{-}}))=\mathbb{U}_{H}\bigl{(}\operatorname{ch}^{\mathbb{T}}(e_{\delta^{-}})\bigr{)}$. The general statement then follows by Corollary 2.6 and the projection formula. Observe that
$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{-}}^{*}e_{\delta^{-}})$ $=\operatorname{ch}^{\mathbb{T}}\biggl{(}i_{\delta^{-}}^{*}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}\biggr{)}$
$=\operatorname{ch}^{\mathbb{T}}\Biggl{(}\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes M_{\delta^{-}_{i}}\bigr{)}^{\vee}\Biggr{)}=\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\Bigl{(}1-{\rm e}^{z_{\delta^{-}_{i}}-z_{j_{-}}}\Bigr{)}$
@@
$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{+}}^{*}\mathbb{FM}(e_{\delta^{-}}))$ $=\operatorname{ch}^{\mathbb{T}}i_{\delta^{+}}^{*}\biggl{(}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}\biggr{)}$
$=\operatorname{ch}^{\mathbb{T}}\Biggl{(}\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes L_{\delta^{+}_{i}}\bigr{)}^{\vee}\Biggr{)}=\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\Bigl{(}1-{\rm e}^{x_{\delta^{+}_{i}}-z_{j_{-}}}\Bigr{)}.$
@@
$\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{\Bigl{(}1-{\rm e}^{x_{\delta^{+}_{i}}-z_{j_{-}}}\Bigr{)}}{\Bigl{(}1-{\rm e}^{z_{\delta^{-}_{i}}-z_{j_{-}}}\Bigr{)}}$
$\qquad{}=\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{{\rm e}^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\Bigl{(}{\rm e}^{-(x_{\delta^{+}_{i}}-z_{j_{-}})/2}-{\rm e}^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\Bigr{)}}{{\rm e}^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\Bigl{(}{\rm e}^{-(z_{\delta^{-}_{i}}-z_{j_{-}})/2}-{\rm e}^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\Bigr{)}}=C_{\delta^{-},\delta^{+}}.$
By the localization theorem in cohomology (see [2]), it follows that the diagram commutes. ∎
@@
${\pi_{-}}_{*}\bigl{(}\alpha\cup\operatorname{Td}_{X_{-}/B}\bigr{)}={\pi_{+}}_{*}\bigl{(}\mathbb{U}_{H}(\alpha)\cup\operatorname{Td}_{X_{+}/B}\bigr{)}.$